OpenStudy (anonymous):

The following function shows the relationship between the selling price (s), and profit P(s), in dollars, for a company. P(s) = -20s2 + 1,400s - 12,000 Which statement best describes the intervals where the company's profit increases, decreases, or records a maximum? It is least when the selling price is $35. It is greatest when the selling price is $35. It decreases when the selling price increases from $10 to $35. It increases when the selling price increases from $35 to $100

OpenStudy (anonymous):

@jdoe0001 @dugalde

OpenStudy (dugalde):

It decreases when the selling price increases beyond $35.

OpenStudy (jdoe0001):

hmm how did you get 35 anyway?

OpenStudy (anonymous):

Please explain @dugalde

OpenStudy (jdoe0001):

anyhow... do you know how to get the vertex of a parabolic equation, or quadratic?

OpenStudy (dugalde):

p(s)=-20(s^2-70s+60)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the vertex is 35, 12500

OpenStudy (anonymous):

@jdoe0001 Do you think the right answer is C

OpenStudy (anonymous):

@jdoe0001 are you there?

OpenStudy (jdoe0001):

the quadratic is has a negative leading term coefficient, -20 meaning the parabola opens downward or  so the maximum profit will occur at the vertex of the parabola \(\bf p(s) = -20s^2 + 1,400s - 12,000\implies p(s)=20(-s^2+70s-600) \\ \quad \\ \textit{vertex of a parabola}\\ \quad \\ p(s) = {\color{red}{ -1}}s^2{\color{blue}{ +70}}s{\color{green}{ -600}}\ \quad \left(-\cfrac{{\color{blue}{ b}}}{2{\color{red}{ a}}}\quad ,\quad {\color{green}{ c}}-\cfrac{{\color{blue}{ b}}^2}{4{\color{red}{ a}}}\right)\)